OpenStudy (anonymous):

Solve the triangle. A = 32°, a = 19, b = 14

hero (hero):

Use law of sines: \[\frac{\sin A}{a} = \frac{\sin B}{b} = \frac{\sin C}{c}\]

hero (hero):

@melchar ?